# Auto-Vocabulary Semantic Segmentation

### v2

@@ 3.3 Self-Guidance
The output of the clustered BLIP embeddings (see Sect. 3.1) is a 32x32 grid with cluster index assignments. Each element of the grid corresponds to an image patch from the original image, which enables us to create segmentation masks - defined as a union of areas covered by image patches with the same index - essentially for free. For example, Fig. 3 shows the clustered output as a 2-dimensional mask, which succeeds in masking relevant objects to a certain extent. However, the segmentation outputs are constrained to a 32x32 resolution of the BLIP encoder. Upsampling yields unsatisfactory results, since objects appear oversegmented and boundaries remain unsharp. Another challenge is to associate each mask with one class name. In this cluster representation, masked areas are described with sentences rather than single nouns. When such a sentence contains multiple nouns, it is not trivial which noun to select as the predicted class name for that mask. To perform effective and accurate auto-vocabulary semantic segmentation, we leverage BCC’s strength in generating an elaborate set of relevant class names and use this as textual guidance for a pre-trained OVS model capable of producing high-resolution outputs. BCC is model-agnostic, therefore any kind of OVS model that expects an image and a set of class labels can be combined with our method. We focus on X-Decoder[42], which is a popular and well-performing OVS model.
3.4 Custom Open-Vocabulary Evaluator
@@
As discussed in Sect. 2, previous works in OVS have mainly focused on a setting where the target class names are provided by the user. Hence, evaluation is possible by having access to the ground truth of those target class names during evaluation. However, in scenarios where the target class names are discovered rather than pre-specified, as in our framework, there may be a lack of direct alignment between the semantics of these categories and the classes used in the annotations. In zero-guidance segmentation [28], the authors have proposed to align class names based on the cosine-similarity in the latent space of Sentence-BERT [26] or CLIP [25]. In our initial experiments, this yielded unsatisfactory results where, even in obvious cases, the open-vocabulary classes were linked to the incorrect class (e.g. taxi to road instead of car). This led to a decline in performance as measured by segmentation metrics, yet the method showed promising results from a qualitative perspective. Relations between two class names can be complex, such as synonymity, hyponymy, hypernymy, meronymy or holonymy, which exceed the capabilities of a cosine-similarity criterium in the respective latent spaces. As a possible solution, we propose an LLM-based Auto-Vocabulary Evaluator (LAVE). Specifically, we employ the Llama 2 7B Large Language model444www.huggingface.co/meta-llama/Llama-2-7b with the task of mapping predicted auto-vocabulary categories to target categories in the dataset. First, AutoSeg returns a segmentation mask with classes from an auto-vocabulary generated by BCC. Second, LAVE collects all uniquely predicted auto-vocabulary categories in the output and builds a mapper from each predicted category to the most relevant or similar target category. Thirdly, all pixel values in the segmentation mask belonging to each class are updated respectively according to the mapper. Finally, the mean Intersection over Union (mIoU) is computed with the updated segmentation mask. Pseudocode and prompts for LAVE are provided in the supplementary material.
@@
Figure 4: Segmentation with VLMs. Example outputs on PASCAL VOC/Context (top), ADE20K (middle) and Cityscapes (bottom) by feeding plain BLIP embeddings to X-Decoder, directly using BCC embeddings as masks and AutoSeg. Notably, our method segments images in the most comprehensive and accurate manner.
@@ 4.1 Experimental Setup
We evaluate our method on three popular semantic segmentation validation datasets: PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] with 20, 459, 847 and 20 classes respectively. These datasets cover a wide range of difficulty and class diversity. For example, ADE20K is challenging with its many and infrequently appearing classes, while Cityscapes contains many object instances per image. We list dataset statistics in Tab. 2.
Evaluation. For quantitative comparison with previous works, we use mIoU as the main metric. As mentioned in Sec. 3.4, class predictions in the segmentation mask are mapped using LAVE before the mIoU is computed. Furthermore, to study the class-agnostic segmentation performance of different VLMs, we report class-agnostic mean Intersection over Union (cmIoU). The computation of cmIoU involves the following steps. First, pairwise IoU values, denoted as $\text{IoU}(g_{i},p_{j})$ are calculated between all ground-truth segments $g_{i}$ and predicted segments $p_{j}$ from a given image. Each $g_{i}$ segment is matched with the predicted segment $p_{\text{max}_{i}}$ that yields the highest IoU value: $p_{\text{max}_{i}}=\arg\max_{j}\text{IoU}(g_{i},p_{j})$. The cmIoU is then determined by calculating the arithmetic mean of these highest IoU values across the dataset, which can be formulated as $\text{cmIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{\text{max}_{i}}$ where $N$ is the total number of ground-truth segments in the dataset.
Table 1: Ablations on Segmentation with VLMs. Class-agnostic mIoU values for different variants of our method. AutoSeg, which combines BCC with X-Decoder segmentation, achieves the best performance on all four datasets.
Approach VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
X-Decoder [42] + BLIP [16] 38.0 35.3 26.7 29.2
Blip-Cluster-Caption 16.3 16.3 11.3 0.85
AutoSeg 71.8 47.7 29.2 35.8
Implementation details. For our experiments, we use the BLIP model built with ViT-Large backbone and finetuned for image captioning on the COCO dataset in combination with the Focal-L variant of X-Decoder. For both models, we use publicly available pretrained weights and do not perform any additional finetuning. Parameter tuning is performed for the clustering and captioning modules. Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 to 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256). As for text generation, we opted for nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible. These parameters were determined using Bayesian optimization on VOC with the goal of maximizing the mIoU.
@@ 4.2 Ablations
To investigate the performance of different configurations of our framework, we present two ablation studies. The first ablation study investigates different methods of using a VLM to obtain segmentation masks. In the second, we examine the effect of different amounts of captioning cycles.
Segmentation with Vision Language Models. A VLM can be leveraged in different ways to obtain segmentation masks. In this ablation study, we study the benefit of using an auto-vocabulary generated with BCC for segmentation as in AutoSeg. In order to do so, we compare our method against a scenario where plain BLIP embeddings are used for captioning instead of enhanced ones. Furthermore, we also evaluate the extent to which BCC embedding outputs themselves can be used directly as segmentation masks (see Sect. 3.3) after upsampling. Due to the difficulty of establishing correspondence between BCC clusters and ground-truth labels, we evaluate the segmentation quality using cmIoU. Quantitative results are reported in Tab. 1. On all four datasets, AutoSeg performs most effectively. Comparing X-Decoder + BLIP to AutoSeg qualitatively, one can observe that the former suffers from false positives and false negatives simultaneously, which in turn serves as an explanation for the performance gap in cmIoU. For example, in PASCAL, it incorrectly masks the area around the bus at the cost of segmenting people. In ADE20K and Cityscapes, it ignores large parts of the area which could appear to BLIP as insignificant. Our framework ensures that smaller, less salient objects are effectively captioned. As expected, intermediate masks of BCC do not segment objects sufficiently accurate due to the low resolution of the embedding space. Interestingly, it is still capable of segmenting global areas linked to certain objects in PASCAL and ADE20K, while failing completely for Cityscapes, likely due to the higher instance count and higher resolution of images.
Table 2: Ablations on Captioning Cycles. The optimal number of captioning cycles varies per dataset. On Cityscapes, which has images in the highest resolution, most instances ($\overline{I}$) and classes on average ($\overline{C}$), AutoSeg benefits from 25 cycles, while between 1 and 15 captioning cycles for datasets with fewer instances per image suffices.
                                 Number of Captioning Cycles
Dataset $\overline{I}$ $\overline{C}$ $C$ 1 5 10 15 20 25
PASCAL VOC [9] 3.5 3.5 20 85.3 84.4 87.1 85.7 83.3 83.3
PASCAL Context [23] 18.9 6.2 459 10.1 9.7 11.2 11.7 9.6 10.8
ADE20K [41] 19.5 10.5 847 5.9 5.8 5.1 5.2 5.1 5.6
Cityscapes [6] 34.3 17 20 28.2 28.1 28.4 28.3 27.9 29.5
@@
Captioning Cycles. A key aspect of our model is its capability to iteratively repeat the captioning process within BCC, enabling it to generate a more exhaustive collection of nouns which accurately describe the scene. This approach proves especially beneficial in instances where objects are initially overlooked or described with less precise semantic terms. To assess the impact, we explore the effects of single versus multiple captioning iterations — 5, 10, 15, 20, or 25 times — using the mIoU metric. The findings, presented in Tab. 2, reveal that the optimal number of captioning cycles varies significantly with the dataset in question. Specifically, we observe that for the Cityscapes dataset, characterized by its high average number of instances and unique classes per image (denoted as $\overline{I}$ and $\overline{C}$ in Tab. 2, respectively), a greater number of cycles yields the most benefit. For datasets with a lower average number of instances and unique classes per image, like VOC or ADE20K, fewer captioning cycles are required for BCC to identify all relevant class names. As anticipated, our model requires a greater number of captioning cycles for the Context dataset, which, despite featuring images identical to those in VOC, includes annotations for an additional 439 object categories. Remarkably, AutoSeg successfully handles ADE20K, the dataset with the largest number of unique classes in our experiments, using merely a single captioning cycle. This outcome underscores our model’s ability in efficiently addressing complex datasets which resemble an open-ended setting.
Table 3: Quantitative Results (mIoU). AutoSeg performs competitive with existing OVS methods which require class names, while achieving superior performance on the zero label setting. Results with dashes are not published and unavailable.
Method Auto-Classes VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
@@
OVSegmentor [36] 53.8 20.4 - -
SimSeg [39] 57.4 26.2 - -
OVDiff [11] 69.0 31.4 - -
OpenSeg [10] 72.2 45.9 8.8 -
OVSeg [17] 94.5 55.7 9.0 -
ODISE [34] 82.7 57.3 11.0 -
SAN [37] 94.6 57.7 12.4 -
CAT-Seg [5] 97.2 62.0 13.3 -
ZSSeg [38] 83.5 47.7 7.0 34.5
X-Decoder [42] 96.2 65.1 6.4 50.8
FC-CLIP [40] 95.4 58.4 14.8 56.2
@@
AutoSeg (Ours) 87.1 11.7 5.9 29.5
4.3 Quantitative Analysis
We compare our method against existing OVS methods that require class names to be given, as well as with Zero-Guidance Segmentation (referred to as ZeroSeg [28]). Tab. 3 shows the results. Without any specification of class names through user input, AutoSeg achieves 86%, 18%, 40% and 52% of the best performing OVS methods on VOC, Context, ADE20K and Cityscapes respectively. The difference in performance can be attributed to AutoSeg needing to identify the subjects for segmentation independently, while OVS methods are explicitly instructed on the specific objects to segment. Remarkably, however, AutoSeg is able to remarkably surpass the performance many OVS methods on VOC, such as OVDiff [11], OpenSeg [10] or ODISE [34], without any access to the ground truth class names. While still leaving room for improvement, our model compares competitively with OVS methods on ADE20K and Cityscapes, two very challenging datasets either high in number of unique classes or instances. Comparing to ZeroSeg, AutoSeg achieves superior performance on VOC (87.1 over 20.1 mIoU) and Context (11.7 over 11.4 mIoU). Furthermore, we set a new state-of-the-art on ADE20K and Cityscapes under the auto-class generation task setting. This outcome underscores the efficacy of our method in dealing with complex scenes that have an open-ended nature, such as ADE20K with its large number of rare classes. While other models have to decide between 847 possibilities, AutoSeg suggests a smaller yet more relevant set of class names. Our method deals well with high numbers of instances and can successfully identify various non-salient objects. A comparison with the findings from Tab. 2 reveals that while BCC struggles with providing high-quality masks on its own, it is able to identify the accurate target classes, which can subsequently be segmented with precision. Additional experiments are provided in the supplementary material, in which we generate class names with other image captioning methods and compare the segmentation performance of such vocabularies. Overall, our results demonstrate that the integration of BCC and OVS harnesses the strengths of both, leading to enhanced open-ended recognition capabilities.
Figure 5: Qualitative Results. AutoSeg shows remarkable capability to identify out-of-vocabulary categories, such as hawk or coke, and segment them accurately across different datasets. The first two rows highlight results on from PASCAL VOC/Context, while the bottom two rows are from ADE20K and Cityscapes respectively.

### v3

@@ 3.2 Segmentation through Self-Guidance
The output of the clustered BLIP embeddings is a 32x32 grid with cluster index assignments (see Sect. 3.1), with each element corresponding to an image patch from the original image. This enables the extraction of segmentation masks - defined as a union of areas covered by image patches with the same index - essentially for free. For instance, Fig.3 and 4 show the clustered output as a 2D mask that partially captures relevant objects. However, the low resolution of the segmentations are unsatisfactory, and upsampling leads to oversegmented objects or unsharp boundaries. To perform effective and accurate auto-vocabulary semantic segmentation, we leverage BBoost’s strength in generating an elaborate set of relevant class names from clusters instead - and use this as textual guidance for a pre-trained OVS model capable of producing high-resolution outputs. BBoost is model-agnostic, allowing our approach to integrate with any OVS model that accepts an image and a set of class labels. In this work, we focus on X-Decoder [48], a popular and well-performing OVS model.
4 Evaluation of Auto-Classes
@@
As discussed in Sect. 2, previous works in OVS have mainly focused on a setting where the target class names are provided by the user. Hence, evaluation is possible by having access to the ground truth of those target class names during evaluation. However, in scenarios where the target class names are discovered, as in our framework, rather than pre-specified there may be a lack of direct alignment between the semantics of these categories and the classes used in the annotations. In zero-guidance segmentation [32], the authors have proposed to align class names based on the cosine similarity in the latent space of Sentence-BERT [30] or CLIP [29]. In our initial tests, however, this approach misaligned obvious class pairs (e.g., taxi mapped to road instead of car), thereby reducing segmentation accuracy despite promising qualitative results. Relations between two class names can be complex and ambigious [44], such as due to synonymity, hyponymy, or hypernymy, which exceed the capabilities of a cosine-similarity criterium in the latent space. To address this, we propose an LLM-based Auto-Vocabulary Evaluator (LAVE), leveraging the Llama-2-7B language model [36], to map predicted auto-vocabulary categories to target dataset classes. LAVE gathers all predicted auto-vocabulary classes and maps each category to the most relevant or similar class in the known vocabulary. It then updates all pixel values in the predicted segmentation masks according to the mapping, after which the mean Intersection over Union (mIoU) is calculated using the updated mask. Though not an integral part of our method, LAVE greatly reduces the mapping effort which is often infeasible to do manually, requiring $C\times A$ comparisons between $C$ known classes and $A$ auto-classes. Pseudocode and prompts for LAVE are included in the supplementary material.
@@
Figure 4: Segmentation with VLMs. Example outputs on PASCAL VOC/Context (top), ADE (middle) and Cityscapes (bottom) by (left to right) directly using BBoost embeddings as masks, feeding plain BLIP embeddings to X-Decoder or AutoSeg. Notably, our method segments images in the most comprehensive and semantically accurate manner.
Table 1: Ablations on Segmentation with VLMs (c-mIoU). Leveraging BBoost semantic embeddings for segmentation with AutoSeg outperforms baselines on all four datasets.
Method VOC [9] PC [27] ADE [47] CS [6]
BBoost Embeddings 16.3 16.3 11.3 0.85
X-Decoder [48] + BLIP [18] 38.0 35.3 26.7 29.2
AutoSeg (Ours) 71.8 47.7 29.2 35.8
@@ 5.1 Experimental Setup
Datasets. AutoSeg is evaluated on four popular semantic segmentation validation datasets: PASCAL VOC [9] and Context (PC) [27], ADE20K (ADE) [47] and Cityscapes (CS) [6] with 20, 459, 847 and 20 classes respectively, covering a wide range of difficulty and class diversity. For instance, ADE is challenging with its many and infrequently appearing classes, while CS contains many instances per image (see Tab. 2 for detailed statistics).
Evaluation. For quantitative comparison with previous works, we use mIoU as the main metric. As mentioned in Sec. 4, auto-class predictions in the segmentation mask are mapped using LAVE before the mIoU is computed. Furthermore, to study the class-agnostic segmentation performance of different VLMs, we report class-agnostic mean Intersection over Union (cmIoU). The computation of cmIoU involves the following steps. First, pairwise IoU values, denoted as $\text{IoU}(g_{i},p_{j})$ are calculated between all ground-truth segments $g_{i}$ and predicted segments $p_{j}$ from a given image. Each $g_{i}$ segment is matched with the predicted segment $p_{\text{max}_{i}}$ that yields the highest IoU value: $p_{\text{max}_{i}}=\arg\max_{j}\text{IoU}(g_{i},p_{j})$. The cmIoU is then determined by calculating the arithmetic mean of these highest IoU values across the dataset, which can be formulated as $\text{cmIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{\text{max}_{i}}$ where $N$ is the total number of ground-truth segments in the dataset.
Implementation details. For our experiments, we use the BLIP model built with ViT-Large backbone and finetuned for image captioning on the COCO dataset in combination with the Focal-L variant of X-Decoder. For both models, we use publicly available pretrained weights and do not perform any additional finetuning. Parameter tuning is performed for the clustering and captioning modules. Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 to 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256). For text generation, we use nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible. These parameters were determined using Bayesian optimization on VOC with the goal of maximizing the mIoU. For CRF denoising and part-of-speech tagging, we use the crfseg [14] and spaCy [11] libraries respectively.
@@ 5.2 Ablations
To assess our framework we investigate various segmentation methods with captions, the impact of captioning cycles, generated/fixed vocabulary similarity, alternative mappers and the effects of cluster denoising. The latter three ablation studies are detailed in the supplementary material.
Segmentation with Vision Language Models. Upsampled BBoost embeddings could potentially act as segmentation masks on their own given the clustering of semantic areas (see Sec. 3.2). In this ablation study, we compare this baseline to AutoSeg which utilizes BBoost’s low-resolution semantics for high-resolution outputs. We assess segmentation quality using class-agnostic (c-mIoU), with results presented in Tab. 1 and Fig. 4. Across all datasets, AutoSeg performs best, with the baseline X-Decoder + BLIP showing issues of false positives and negatives that explain the performance gap. For example, it incorrectly segments areas around objects in VOC/PC (e.g. person for building) and ignores significant regions in ADE and CS. In contrast, AutoSeg effectively captures smaller objects and details. As expected, low-resolution BBoost masks are inaccurate but can still segment general areas in VOC/PC and ADE, though they struggle in CS due to higher object density.
Table 2: Ablations on Captioning Cycles (mIoU) and Generated Classes. The optimal number of captioning cycles depends on the dataset, influenced by the image content and annotation method. Our method identifies a substantially larger variety of unique classes (Gen.) than those annotated by humans (Ann.), highlighting its capacity to capture finer context. $\overline{I}$ and $\overline{C}$ denote the average number of instances and classes per image.
Number of Captioning Cycles Data Properties
Dataset 1 5 10 15 20 25 Gen. Ann. $\overline{I}$ $\overline{C}$
PASCAL VOC [9] 85.3 84.4 87.1 85.7 83.3 83.3 938 20 3.5 3.5
PASCAL Context [27] 10.1 9.7 11.2 11.7 9.6 10.8 721 459 18.9 6.2
ADE20K [47] 6.0 5.8 5.1 5.2 5.1 5.6 1578 847 19.5 10.5
Cityscapes [6] 28.2 28.1 28.4 28.3 27.9 30.0 395 20 34.3 17
@@
Captioning Cycles. A key aspect of our model is its capability to repeatedly enrich the vocabulary with BBoost, enabling it to generate a more exhaustive collection of classes which accurately describe the scene. This approach proves especially beneficial for instances where objects are initially overlooked or described with less precise semantic terms. To assess its impact, we explore the effects of various captioning iterations using the mIoU metric. The findings, presented in Tab. 2, reveal that the optimal number of captioning cycles varies significantly with the dataset in question. Specifically, we observe that for CS, characterized by its high number of instances and unique classes per image (denoted as $\overline{I}$ and $\overline{C}$ in Tab. 2, respectively), a greater number of cycles yields the most benefit. For datasets with a lower average number of instances and unique classes per image, like VOC or ADE, fewer captioning cycles are required for BBoost to identify all relevant class names. As anticipated, our model requires a greater number of captioning cycles for the CS dataset given its additional 439 object categories. Remarkably, AutoSeg successfully handles ADE, the dataset with the largest number of unique classes in our experiments, using merely a single captioning cycle. This outcome underscores our model’s ability in efficiently addressing complex datasets which resemble an open-ended setting. Finally, we note that through its captioning framework, AutoSeg perceives significantly more distinct classes compared to the hand-crafted fixed vocabularies.
5.3 Quantitative Analysis
Auto-Vocabulary Segmentation Setting. Our quantitative analysis begins by comparing AutoSeg with other auto-vocabulary methods, such as Zero-Guidance Segmentation (ZeroSeg [32]), as well as LLM-aided segmentation methods like LISA [16] with alternative captioning methods like LLaVA [21]. We also evaluate a configuration combining BLIP and X-Decoder, where instance crops predicted by SAM [13] are used for captioning to simulate locality. We use the same caption filtering as in our method. The results, shown in Tab. 3, indicate that our method outperforms ZeroSeg, the only other true auto-vocabulary approach. In contrast, LISA requires multiple inferences with individually prompted categories to achieve scene segmentation—a notably more complex and less efficient approach that also yields poor performance. Captioning based on smaller object crops is similarly ineffective compared to using enhanced vision-language features that maintain spatial locality. Moreover, while LISA provides LLM-guided captions, these lack the contextual specificity needed to establish an effective vocabulary for X-Decoder. Although AutoSeg also utilizes this segmentation backbone, it achieves a substantially higher performance. This comparison not only highlights the unique strengths of AutoSeg but also reveals the practical limitations of current LLM-aided approaches in segmenting scenes with multiple object categories.
Table 3: Quantitative Auto-Vocabulary Results (mIoU). The wide availability of captioning and OVS methods allows various combinations to design an auto-vocabulary method, but they are not necessarily performing well. AutoSeg performs superior over the only Auto-Vocabulary method ZeroSeg, as well as over AVS-adapted methods with alternative captioners or segmentors.
Auto-Vocabulary Segmentation Method VOC [9] PC [27] CS [6]
LLaVA + LISA [21, 16] 7.7 0.2 1.5
ZeroSeg [32] 20.1 11.4 -
SAM + BLIP + X-Decoder [13, 18, 48] 41.1 11.3 27.4
LLaVA + X-Decoder [21, 48] 56.7 11.4 23.4
AutoSeg (Ours) 87.1 11.7 30.0
Figure 5: Qualitative Results. AutoSeg shows remarkable capability to identify out-of-vocabulary categories, such as hawk or coke, and segment them accurately across different datasets. Images are from the VOC/PC, ADE and CS datasets.
Table 4: Open/Auto-Vocabulary State of the Art Comparison (mIoU). AutoSeg surpasses ZeroSeg [32] in human-free segmentation and remains competitive with some well-known OVS methods that rely on human input, while outperforming others (3 out of 8 methods on VOC and 3 out of 7 methods on PC). Results with dashes indicate unpublished or unavailable data.
Unknown VOC [9] PC [27] ADE [47] CS [6]
Method Vocabulary (20) (459) (847) (20)
OVS Segmentation methods with prompted (known = ground truth) vocabulary
@@ 5.3 Quantitative Analysis
OpenSeg [10] 72.2 9.0 8.8 -
OVSeg [19] 94.5 11.0 9.0 -
ODISE [41] 82.7 13.8 11.0 -
SAN [43] 94.6 12.6 12.4 -
CAT-Seg [5] 97.2 19.0 13.3 -
X-Decoder [48] 96.2 16.1 6.4 50.8
FC-CLIP [46] 95.4 12.8 14.8 56.2
AVS Segmentation methods with auto-generated (unknown) vocabulary
@@
AutoSeg + LAVE mapper (Ours) 87.1 11.7 6.0 30.0
AutoSeg + Manual mapper (Ours) 88.2 12.8 6.2 31.1
Open-Vocabulary Segmentation Setting. We compare our method against existing OVS methods that require class names to be given. Tab. 4 shows the results. In addition to the results mapped with LAVE, we provide results with one manual mapper per dataset (note that we provide a manual mapper for this table only, given the significant manual effort to construct it). Without any specification of class names through user input, AutoSeg matches 91%, 67%, 42% and 55% of the best OVS method performance on VOC, PC, ADE and CS respectively. It should be noted that this metric reflects the performance on the known, annotated classes, while additional open-ended classes are potentially mapped. Despite not being explicitly instructed with the known vocabulary contrary to OVS methods, AutoSeg is remarkably able to surpass six of them on VOC and PC, such as OpenSeg [10], ODISE [41] or OVSeg [19]. While still leaving room for improvement, our model compares competitively with OVS methods on ADE and CS, two very challenging datasets either high in number of unique classes or instances. Compared to ZeroSeg, AutoSeg achieves superior performance on VOC (88.2 over 20.1 mIoU) and PC (12.8 over 11.4 mIoU). Furthermore, we set the first benchmark on ADE and CS under the unknown vocabulary task setting. This outcome underscores the efficacy of our method in dealing with complex scenes that are open-ended in nature, such as ADE with its large number of rare classes. Finally, results obtained with LAVE mappings show little difference with manually constructed ones. This indicates that at marginal cost, LAVE can act as a feasible bridge between a known and unknown vocabulary. Our method deals well with high numbers of instances and can successfully identify various non-salient objects. While BBoost struggles with producing high-quality masks by itself (as seen in Tab. 2), it is highly effective in embedding target classes accurately which can be segmented with precision afterwards. Overall, our results demonstrate that the integration of BBoost with OVS harnesses the strengths of both, leading to enhanced open-ended recognition capabilities.